OpenStudy (anonymous):

True or False: (-101 + x) + (101 - x) = 0 is this true right?

OpenStudy (callisto):

Yes, (-101 + x) + (101 - x) = -101 +101 +x-x = 0 :)

OpenStudy (anonymous):

YAY

OpenStudy (callisto):

congratulations :D